# Superconducting spintronics with electron symmetry filtering and interfacial spin-orbit coupling

### v1

@@ IV.3 Conductance anomalies: disentangling SOC and spin texture contributions to LRTs formation
Figure 14: CAs in a F/S/F junction. (a) Conductance vs bias below (blue), near (green), and above (red) $T_{C}$, showing how the CAs disappear in the normal state. (b) CA amplitude near 15-15 mV vs $T$ for different applied fields, IP and OOP. The lines are guides for the eyes, and the arrows indicate the estimated $T_{C}$ of each experiment. The inset shows the conductance around the CA below and above $T_{C}$. (c) Conductance vs bias of the same junction at $T=0.3$ K for different IP applied fields, showing the robustness of the CAs under large magnetic fields. (d) amplitude of a CA vs applied field, both IP and OOP. The inset shows an IP MR measured at $T=0.3$ K, showing that the IP magnetization does not fully saturate until an applied field of about 0.3 T. Adapted from Ref.Tuero2024 with the necessary permissions.
@@
Similar CAs were also observed in F/S/F junctions (the structure of these samples is sketched in Figure 1d, and the data is shown in Figure 14a). They are also robust under high applied magnetic fields, and show a sharp transition near $T_{C}$ (Figure 14b,c). However, as shows Figure 14d, the behavior of CAs on IP and OOP fields is in some way different to one observed in S/F/F junctions. While in both types of junctions the CAs maintain a baseline level in the high field regime, in the F/S/F samples the CA enhancement at low fields is seen both for IP and OOP fields. This is the result of the F/S/F junctions having a much stronger AP coupling compared to the S/F/F ones, which could induce additional STs in the low IP field regime. Combined with the PMA, this results in the magnetization only becoming fully saturated under an IP applied field above $0.3$ T (see Figure 14d inset, showing an IP magnetoresistance). In contrast, this extra amplitude of the CAs disappears more quickly with an OOP applied field, as in this case the PMA and the applied field point in the same direction. The CAs still persist in the high field regime (Figure 14c,d) supporting a scenario of dominant SOC-induced LRT pair generation with an additional contribution of ST-LRTs in the low field regime.
@@ V Excess subgap shot noise in S/F junctions
Another posible explanation could be based in previous results Wei2010, which showed that strong low-bias conductance anomalies could result in corresponding anomalies in shot noise signal. If this was the case, however, we used the same model to estimate that the shot noise signal in our junctions could increase up to 1.4 times, while the experiments show a 1010010-100 times increase.
The next step is to turn our attention to intrinsic phenomena: giant shot noise was found in S/S Dieleman1997 and S/N/S Hoss2000 junctions and explained with multiple Andreev reflections (MAR) Averin1996; Cuevas1999. The advantage of using MAR to expalain the observed shot noise resides in that (due to bunching) MAR transfers a larger effective charge, which can result in huge subgap shot noise at low biases. However, to the best of our knowledge, such effect has not been reported so far in devices with only a single superconducting layer. Taking this into account, the model used is based on the possibility of the inducing (through proximity effect) of a second interfacial superconducting region in the junction inside the F layer GonzalezRuano2025. Then, SOC could allow spin-flip Andreev reflections to form LRTs. This is basically what happens in Josephson junctions (JJs) Octavio1983; Dieleman1997, where noise excess well above the Poissonian levels have been reported and explained by multiple Andreev reflections. A proximity-induced gap in a ferromagnet near a superconductor was also predicted in previous works Huertas-Hernando2005; Golubov2012. Similar phenomena was also suggested to explain experimental data resembling Josephson effects in S/N structures Han1990. Figure 15d sketches our proposed model. In the model, ehe-h quasi-particles with $\Delta_{1}$ symmetry are available near the Fermi level of the Fe electrode, and could be transferred through the MgO Butler2001. The ehe-h pairs would then be reflected multiple times before finally penetrating in the opposite electrode, producing the huge increase of the shot noise signal in the superconducting regime.
@@
VI Outlook: Josephson Junctions based on S/SOC/F systems
@@
Figure 16: (a) SEM image of the lateral junctions under study with the zone attacked by FIB in the center. The yellow arrows and the angle $\theta$ define the coordinate system used for the external magnetic field. The white arrow defines the sign of the current with respect to the coordinate system. In the inset, a sketch of vertical structure of the junctions under study and the electronic set-up for the measurements. (b) Conductance versus temperature $T$ with $\mu_{0}H=0$ measured at two different biases. (c) Differential conductance versus bias curves at $\mu_{0}H=0$ for different temperatures. (d) Conductance versus angle of the applied IP external magnetic field $\theta$ for different current densities across the junction, that correspond with biases of about 1.20 mV, 0.54 mV and 0.31 mV, from highest to lowest. (e) Conductance versus temperature curves for different $\theta$. In the inset, conductance at $T=0.3$ K as a function of $\theta$.
@@
To sum up this section, we have presented promising results which point towards the possible presence of LRTs in our lateral Josephson junctions. We do not completely exclude the possible presence of pinholes in these junctions. Further detailed research is planned to fully understand the above-described phenomena and gather conclusive evidence on the possibility of controlling spin supercurrents via the magnetization orientation in lateral S/F/S JJs with interfacial SOC.
VII Summary and conclusions
@@
In this work, we have focused on the anomalous electron transport in SF junctions with SOC, leaving aside several other important effects that have recently emerged in S/F hybrids with SOC as they are not directly related to epitaxy and, therefore, orbital symmetry filtering. One such effect is ferromagnetic resonance (FMR)-induced spin supercurrent pumping Yao2018; Jeon2020. Our currently ongoing dynamic experiments aim to investigate the temperature dependence of damping in V/MgO/Fe-based junctions. Preliminary results suggest an enhanced FMR damping in the superconducting state compared to the normal state, which could provide independent evidence of LRT generation induced by SOC localized at the S/F interface (in contrast to previous reports where SOC was localized aside the S and F layers Jeon2020). Additionally, we have not discussed the magnetic control of giant thermoelectric effects in V/MgO/Fe based junctions ouassou-prb-22; GonzlezRuano2023; Tuero2025, since both SOC and symmetry filtering are not critical for their observation. Recent theoretical work, however, predicts that the symmetry of the thermoelectric response in S/F junctions with SOC could serve as yet another probe of LRT generation Dutta2020. Finally, we have not discussed LRTs generation in superconductor-antiferromagnet hybrids Jeon2021, where the role of SOC (without involving orbital symmetry filtering) has only recently been addressed by theory Jakobsen2020; Johnsen2021; Bobkov2023.
@@ Acknowledgements
We thank Diego Caso, Isidoro Martínez and Juan Pedro Cascales for their help with experiments and simulations. Jaroslav Fabian, Jacob Linder, Jong Han, Chenghao Shen, Michel Hehn, Pilar Prieto, Andreas Costa and Alex Matos-Abiague are also acknowledged for the fruitful discussions and collaboration. Finally, Manuel Rodríguez and Fernando Jiménez from IMDEA Nanoscience are acknowledged for their help with fabrication of lateral Josephson junctions. The work in Madrid was supported by Spanish Ministry of Science and Innovation (PID2021-124585NB-C32, TED2021-130196B-C22 and PID2024-155399NB-I00). F.G.A. also acknowledges financial support from the Spanish Ministry of Science and Innovation through the María de Maeztu Programme for Units of Excellence in R&D (CEX2023-001316-M). This research also received support from the Comunidad de Madrid through project TEC-2024/TEC-380 (Mag4TIC-CM). The growth of samples was performed using equipment from the CC-DAUM platform funded by FEDER (EU),ANR, the Region Lorraine, and the metropole of Grand Nancy. I.Ž. acknowledges U.S. DOE, Office of Science BES, Award No. DE-SC0004890.

### v2

@@ IV.3 Conductance anomalies: disentangling SOC and spin texture contributions to LRTs formation
Figure 14: Conductance anomalies (CAs) in a F/S/F junction. (a) Conductance vs bias below (blue), near (green), and above (red) $T_{C}$, showing how the CAs disappear in the normal state. (b) CA amplitude near 15-15 mV vs $T$ for different applied fields, in-plane (IP) and out-of-plane (OOP). The lines are guides for the eyes, and the arrows indicate the estimated $T_{C}$ of each experiment. The inset shows the conductance around the CA below and above $T_{C}$. (c) Conductance vs bias of the same junction at $T=0.3$ K for different IP applied fields, showing the robustness of the CAs under large magnetic fields. (d) amplitude of a CA vs applied field, both IP and OOP. The inset shows an IP magnetoresistance measured at $T=0.3$ K, showing that the IP magnetization does not fully saturate until an applied field of about 0.3 T. Adapted from Ref.Tuero2024 with the necessary permissions.
@@
Similar CAs were also observed in F/S/F junctions (the structure of these samples is sketched in Figure 1d, and the data is shown in Figure 14a). They are also robust under high applied magnetic fields, and show a sharp transition near $T_{C}$ (Figure 14b,c). However, as shown in Figure 14d, the behavior of CAs with IP and OOP fields is in some way different to one observed in S/F/F junctions. While in both types of junctions the CAs maintain a baseline level in the high field regime, in the F/S/F samples the CA enhancement at low fields is seen both for IP and OOP fields. This is the result of the F/S/F junctions having a much stronger AP coupling compared to the S/F/F ones, which could induce additional STs in the low IP field regime. Combined with the PMA, this results in the magnetization only becoming fully saturated under an IP applied field above $0.3$ T (see Figure 14d inset, showing an IP magnetoresistance). In contrast, this extra amplitude of the CAs disappears more quickly with an OOP applied field, as in this case the PMA and the applied field point in the same direction. The CAs still persist in the high field regime (Figure 14c,d) supporting a scenario of dominant SOC-induced LRT pair generation with an additional contribution of ST-LRTs in the low field regime.
@@ V Excess subgap shot noise in S/F junctions
Another posible explanation could be based in previous results Wei2010, which showed that strong low-bias conductance anomalies could result in corresponding anomalies in shot noise signal. If this were the case, our model predicts that the shot noise signal in our junctions would increase by at most a factor of 1.4, whereas the experiments show a 1010010-100 times increase.
The next step is to turn our attention to an intrinsisc phenomenon: giant shot noise was found in S/S Dieleman1997 and S/N/S Hoss2000 junctions and explained with multiple Andreev reflections (MAR) Averin1996; Cuevas1999. The advantage of using MAR to expalain the observed shot noise resides in that (due to bunching) MAR transfers a larger effective charge, which can result in huge subgap shot noise at low biases. However, to the best of our knowledge, such effect has not been reported so far in devices with only a single superconducting layer. Taking this into account, the model used is based on the possibility of the inducing (through proximity effect) of a second interfacial superconducting region in the junction inside the F layer GonzalezRuano2025. Then, SOC could allow spin-flip Andreev reflections to form LRTs. This is basically what happens in Josephson junctions (JJs) Octavio1983; Dieleman1997, where noise excess well above the Poissonian levels have been reported and explained by multiple Andreev reflections. A proximity-induced gap in a ferromagnet near a superconductor was also predicted in previous works Huertas-Hernando2005; Golubov2012. Similar phenomena was also suggested to explain experimental data resembling Josephson effects in S/N structures Han1990. Figure 15d sketches our proposed model. In the model, ehe-h quasi-particles with $\Delta_{1}$ symmetry are available near the Fermi level of the Fe electrode, and could be transferred through the MgO Butler2001. The ehe-h pairs would then be reflected multiple times before finally penetrating in the opposite electrode, producing the huge increase of the shot noise signal in the superconducting regime.
@@
VI Josephson Junctions based on S/SOC/F systems
@@
Figure 16: (a) SEM image of the lateral junctions under study with the zone attacked by focus ion beam (FIB) in the center. The yellow arrows and the angle $\theta$ define the coordinate system used for the external magnetic field. The white arrow defines the sign of the current with respect to the coordinate system. In the inset, a sketch of vertical structure of the junctions under study and the electronic set-up for the measurements. (b) Conductance versus temperature $T$ with $\mu_{0}H=0$ measured at two different biases. (c) Differential conductance versus bias curves at $\mu_{0}H=0$ for different temperatures. (d) Conductance versus angle of the applied in-plane external magnetic field $\theta$ for different current densities across the junction, that correspond with biases of about 1.20 mV, 0.54 mV and 0.31 mV, from highest to lowest. (e) Conductance versus temperature curves for different $\theta$. In the inset, conductance at $T=0.3$ K as a function of $\theta$.
@@
To sum up this section, we have presented promising results which point towards the possible presence of LRTs in our lateral Josephson junctions. We do not completely exclude the possible presence of pinholes in these junctions. Further detailed research is planned to fully understand the above-described phenomena and gather conclusive evidence on the possibility of controlling spin supercurrents via the magnetization orientation in lateral S/F/S JJs with interfacial SOC. Interestingly, theoretical predictions reveal a nonmonotonic spin-triplet contribution with the strength of the interfacial barrier and spin-orbit coupling Shen2024.
VII Outlook
The structure of this manuscript reflects the fact that Fe/MgO/V-based heterostructures are currently the only systems in which superconducting spintronics involves the interplay of Cooper pairing, electron symmetry filtering, and interfacial spin-orbit coupling. Similar studies would be of interest in superconductor-antiferromagnet (S/AF) juntions with SOC. Indeed, recently, attention has turned to S/AF hybrids. Although most of this works are theoretical, they suggest that S/AF systems, despite having nearly zero net magnetization, could also host non-trivial spin-triplet correlations when SOC or non-collinear spin textures are present. Notably, the mechanisms generating long-range triplet Cooper pairs in these systems differ from those in conventional S/F structures, as they do not rely on exchange splitting of otherwise degenerate energy levels. Jakobsen and co-workers Jakobsen2020 theoretically examined electrical and thermal transport in S/AF junctions with Rashba SOC. They predict that rotating the Néel vector from an in-plane orientation to one perpendicular to the interface can produce a large low-bias anisotropic magnetoresistance, qualitatively similar to the magneto-anisotropic Andreev reflection effect discussed earlier for S/F systems with Rashba SOC (see Chapter IV.1 and Ref.Martinez2020).
In agreement with these findings, Bobkov et al. report that Rashba SOC can anisotropically modify proximity-induced spin-triplet correlations, leading to an orientation-dependent modulation of the superconducting critical temperature as the Néel vector is rotated relative to the S/AF interface Bobkov2023. To account for a stronger-than-expected suppression of the critical temperature in S/AF bilayers, another mechanism that does not involve SOC has been proposed. In compensated antiferromagnets, the superconducting proximity effect can generate Néel-triplet Cooper pairs whose pairing amplitude oscillates spatially in step with the AF lattice spin structure Bobkov2022. While this mechanism weakens superconductivity in clean systems, interfacial disorder that induces a local ferromagnetic moment offers an alternative route for generating spin-triplet correlations in S/AF junctions. Fyhn et al. predict that in disordered metallic antiferromagnets, impurities (especially those near the interface) that promote interfacial ferromagnetism, suppress singlet correlations but leave triplet correlations largely unaffected Fyhn2023. Adjusting the direction of this interfacial ferromagnetic moment relative to the bulk Néel vector can selectively filter the spin-triplet component of the proximity effect, providing a potential means to realize a magnetic switch for spin supercurrents in S/AF junctions.
In this work, we have focused on the anomalous electron transport in S/F junctions with SOC, leaving aside several other important effects that have recently emerged in S/F hybrids with SOC as they are not directly related to epitaxy and, therefore, orbital symmetry filtering. One such effect is ferromagnetic resonance (FMR)-induced spin supercurrent pumping Yao2018; Jeon2020. Our currently ongoing dynamic experiments aim to investigate the temperature dependence of damping in V/MgO/Fe-based junctions. Preliminary results suggest an enhanced FMR damping in the superconducting state compared to the normal state, which could provide independent evidence of LRT generation induced by SOC localized at the S/F interface (in contrast to previous reports where SOC was localized aside the S and F layers Jeon2020). Additionally, we have not discussed the magnetic control of giant thermoelectric effects in V/MgO/Fe based junctions ouassou-prb-22; GonzlezRuano2023; Tuero2025, since both SOC and symmetry filtering are not critical for their observation. Recent theoretical work, however, predicts that the symmetry of the thermoelectric response in S/F junctions with SOC could serve as yet another probe of LRT generation Dutta2020.
Finally, we believe that the above discussed effects could be explored in other epitaxial heterostructures such as V/MgO/Co or Nb/MgO/Fe - based systems. In the junctions involving Co instead of Fe, the thicknes of Co layer should be only of few nanometers. This is because ultrathin Co grown on top of MgO could transform its hexagonal crystalline structure to a body-centered one, and provide even better orbital symmetry and spin filtering effects than the one taking place in V/MgO/Fe Yuasa2006. On the other hand, replacing vanadium with bcc Niobium could probably not be feasible for superconducting spintronics with electron symmetry filtering and interfacial spin-orbit coupling. First, the lattice mismatch between superconductor and MgO would increase by roughly a factor of three Guasco2023, thereby degrading the epitaxial growth and enhancing the concentration of extended defects such as dislocations with a significant effect on noise. Secondly, as niobium has pronounced $\Delta_{1}$ states at the Fermi level Mattheiss1970, the impact of SOC on electron and spin transport in Nb/MgO/Fe heterostructures would be greatly reduced due to an additional shunting current channel between the electrodes created by the $\Delta_{1}$ “hot spots” in MgO.
VIII Summary and conclusions
@@
Data availability statement
The data that support the findings of this study are available upon reasonable request from the authors.
@@ Acknowledgements
We thank Diego Caso, Isidoro Martínez and Juan Pedro Cascales for their help with experiments and simulations. Jaroslav Fabian, Jacob Linder, Jong Han, Chenghao Shen, Michel Hehn, Pilar Prieto, Andreas Costa and Alex Matos-Abiague are also acknowledged for the fruitful discussions and collaboration. Finally, Manuel Rodríguez and Fernando Jiménez from IMDEA Nanoscience are acknowledged for their help with fabrication of lateral Josephson junctions. The work in Madrid was supported by Spanish Ministry of Science and Innovation (PID2021-124585NB-C32, TED2021-130196B-C22 and PID2024-155399NB-I00). F.G.A. also acknowledges financial support from the Spanish Ministry of Science and Innovation through the María de Maeztu Programme for Units of Excellence in R&D (CEX2023-001316-M). This research also received support from the Comunidad de Madrid through project TEC-2024/TEC-380 (Mag4TIC-CM). The growth of samples was performed using equipment from the CC-DAUM platform funded by FEDER (EU),ANR, the Region Lorraine, and the metropole of Grand Nancy. I.Ž. acknowledges U.S. DOE, Office of Science BES, Award No. DE-SC0004890. C.T. acknowledges the funding project UEFISCDI via the project “MODESKY” ID PN-III-P4-ID-PCE-2020-0230-P, Grant No. UEFISCDI: PCE 245/02.11.2021.